# Iterative Prompt Relocation for Distribution-Adaptive Visual Prompt Tuning

### v1

@@ 3.3 Iterative Prompt Relocation-based VPT
where the idleness score ^$\widehat{\mathcal{I}}_{k^{\!\ast}}\!\approx\!\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P},\mathcal{D}^{-}})$ provides an approximation of the intermediate loss of $f_{\mathbf{P},\mathcal{D}^{-}}$.
By leveraging RL, we can efficiently determine the optimal allocation block for the pruned prompt, even within the non-stationary environment. We achieve this using proximal policy optimization (PPO) [44], a state-of-the-art RL algorithm. In particular, PPO learns a policy that identifies the optimal block to maximize the reward in Eq. (8).
Input: pre-trained model $f$, number of epochs $T$, number of prompts $N$, learning rate $\eta$.
1 Initialize $N$ prompts $\mathbf{P}$ and distribute them to the model according to a uniform distribution $\mathcal{D}$.
2 for $t=0,\ldots,T-1$ do
3       Compute idleness scores {^$\{\widehat{\mathcal{I}}_{k}\}^{N}_{k=1}$ by Eq. (6).
4       if $\max\widehat{\mathcal{I}}_{k}\!>\!0$ then
5             Prune the negative prompt $\mathbf{p}_{k^{\!\ast\!}}$ with the maximum idleness score ^$\widehat{\mathcal{I}}_{k^{\!\ast\!}}$ to form $\mathcal{D}^{-}$.
6             Compute the current state $s$.
7             Compute the action $a\leftarrow\text{PPO}(s)$.
8             Allocate the idle prompt $\mathbf{p}_{k^{\!\ast\!}}$ to the $a^{\text{th}}$ block to form $\mathcal{D}^{+}$.
9
11       if $\max\widehat{\mathcal{I}}_{k}\!>\!0$ then
12             Compute the reward $\hat{r}$ by Eq. (8).
13             Update the policy networks within PPO based on $\hat{r}$.
14
Algorithm 1 The proposed PRO-VPT algorithm
Justification. Although we decompose the PR process into two-step sequential steps, the overall optimization objective, whether it is based on the expected objectives in Eqs. (3) and (4) or the estimated objectives in Eqs. (6) and (8), can be expressed as $\max_{a}\mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}[\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P}^{\prime},\mathcal{D}\mid d_{k^{\!\ast}\!=\!a}})]$. This formulation remains equivalent to the original PR objective (i.e., the ADO objective) defined in Eq. (1). Consequently, leveraging the nested optimization formulation, the proposed PR strategy can effectively address the ADO problem and adaptively construct the optimal distribution for the entire prompt set.
Overall PRO-VPT Framework. Fig. 3 illustrates the streamlined process of PRO-VPT, and Algorithm 1 presents the detailed pseudocode.
@@ 3.4 Discussion
We note that some prior works have indirectly achieved adaptive distribution [13, 25, 60]. Specifically, pruning-based techniques [13, 25] introduce mask variables to iteratively prune redundant prompts from the original distribution, and NOAH [60] use an evolutionary algorithm to search for an optimal sub-structure within the original distribution in one-shot manner. Overall, these approaches aim to enhance performance through the lottery ticket hypothesis [8, 39], focusing on searching for a sub-distribution within a given hyper-distribution. PRO-VPT is distinct and advantageous compared to these methods as follows:
Distinct Goals. These methods primarily concentrate on circumventing negative prompts and retaining essential ones to enhance performance. In contrast, our approach actively learns an optimal distribution of all prompts, relocating them to maximize the effectiveness of each prompt. This fundamental distinction sets our approach apart from the methods discussed above.
Efficient. While effective, we argue that these approaches typically incur considerable unnecessary training costs, as numerous trained prompts are pruned and left unused (cf. § 4.3). This fundamentally contradicts the goal of parameter efficiency. Conversely, our work directly adjusts the distribution of the entire prompt set through prompt relocation, thereby minimizing extra costs and improving efficiency.
Effective. These methods straightforwardly apply one-shot or iterative processes without considering the underlying distribution dynamics. For example, NOAH adopts a one-shot approach post-training, leading to suboptimal results (cf. § 4.2). Conversely, our work conducts empirical analyzes and proposes a nested-optimization-based method, enabling prompt-based methods to reach state-of-the-art performance.
@@ 4.1 Experiment Setups
Datasets. We mainly evaluated our proposed PRO-VPT on two standard benchmarks for task adaptation. VTAB-1k [59] consists of 19 benchmarked Visual Task Adaptation, categorized into three groups: (1) Natural includes natural images captured by standard cameras; (2) Specialized contains images captured by specialized equipment, from remote sensing and medical domains; (3) Structured covers tasks requiring geometric comprehension such as object counting. Each dataset in VTAB-1k includes 800 training images and 200 validation images, with the test sets matching the number of images in the original datasets. FGVC contains 5 benchmarked Fine-Grained Visual Classification, including CUB-200-2011 [51], NABirds [50], Oxford Flowers [40], Stanford Dogs [6], and Stanford Cars [10]. Since VTAB-1k is designed to assess task adaptation under limited training data conditions, we use FGVC to evaluate adaptation methods in settings where training data is large-scale.
Implementation Details. We mainly use ViT-B/16 [7] supervised pre-trained on ImageNet-21k [43] as the initialization. We employ the SGD optimizer with a batch size of 64 and fine-tune models for 100 epochs. The number of prompt tokens per task in our method is set in accordance with [20]. Following [47], we evaluate our method using Inception normalization instead of ImageNet normalization, and for fair comparison, we present the results of other counterparts using both normalization techniques. Notably, unlike most approaches [61, 9, 20], our method employs a fixed learning rate instead of using a learning rate schedule, as we necessitate ensuring timely rewinding after prompt relocation.
More details about the datasets and the hyperparameters are provided in Appendices F and G.
@@ 4.2 Main Results
We compare the proposed PRO-VPT with the following state-of-the-art methods: (1) Traditional fine-tuning baselines, including full fine-tuning and linear probing; (2) Adapter-based methods, including SSF [31], SPT-Adapter [16], and Adapter+ [47]; (3) Reparameterization-based methods, including LoRA [18], FacT [21], and Consolidator [15]; (4) Prompt-based methods, including VPT [20], NOAH [60], SPT [53], and iVPT [62].
VTAB-1k. We present our results on VTAB-1k with the best-performing normalization techniques (i.e., ImageNet or Inception) for each method, as shown in Tab. 1. Comprehensive results, including both normalization types, are available in Appendix G. Among all prompt-based methods, the previous best performance of 77.1% average accuracy achieved by iVPT falls short of the advanced adapter-based method Adapter+, which reaches 77.6%. Nevertheless, with the introduction of the optimization-based adaptive distribution, the proposed PRO-VPT sets a new state-of-the-art performance among all evaluated methods with an average accuracy of 78.0%, highlighting the significant potential of prompt-based approaches.
@@
Furthermore, Fig. 4 illustrates the performance differences between VPT w/ ADO (PRO-VPT) and w/o ADO (VPT-Deep) on the VTAB-1k datasets. The accuracies of PRO-VPT are consistently higher than those of VPT-Deep, yielding an average accuracy improvement of 1.6%. This emphasizes the significance of optimizing the prompt distribution to unlock the full potential of VPT.
FGVC. Next, we present the FGVC results with the best-performing normalization in Tab. 2. Complete results for both normalizations are also provided in Appendix G. PRO-VPT achieves the best average accuracy of 91.7% over all five FGVC datasets, surpassing the second-best performance of 91.4% achieved by Adapter+ and SPT-Deep. This demonstrates that PRO-VPT also delivers state-of-the-art results for task adaptation when training data is large-scale.
@@ 4.3 Ablation Study
Strategies (A), (B) represent alternatives for extracting a sub-distribution from a hyper-distribution (see § 3.4). We specifically utilize the pruning-based technique as in [13], applying pruning percentages ranging from 10% to 90% with 10% intervals. In particular, (B) prunes from a huge hyper-distribution to obtain a sub-distribution with the same final parameter count as the VPT baseline, and (A) prunes from a smaller hyper-distribution to match the number of trained parameters in the VPT baseline.
Strategies (C)-(G) represent alternatives of prompt relocating, each incorporating different components. Particularly, (C) frames the PR process as a naive RL problem, as discussed in § 3.3, and utilizes the PPO algorithm. (D)-(G) are various combinations of the following components:
@@
Observation 3. Both the pruning and allocation components in PRO-VPT exhibit significantly effective, especially RL-based allocation. Comparing strategies (D) vs. (F) and (E) vs. (G), Prn2 indicates superior performance improvements (averaging 1.07% for SVHN and 0.79% for Resisc45), highlighting the effectiveness of the proposed idleness score. Moreover, strategies (G) and (E) achieve the best results, with Alc2 delivering remarkable gains over Alc1 (2.24% and 1.71%), highlighting the advantage of framing the allocation task as a non-stationary RL problem.
@@ 4.4 Generalizability Analysis
Pre-training Strategy. In addition to the supervised pre-training strategy used in Tab. 1, we experiment with two self-supervised strategies: the masked image modeling method (MAE) [17] and the contrastive self-supervised method (MoCo-v3) [3]. As illustrated in Tab. 5, PRO-VPT consistently outperforms VPT regardless of the pre-training strategy, highlighting its strong generalizability.
Detection and Segmentation. See Appendix G.
@@
Figure 6: Cosine similarity between feature changes and corresponding test accuracy. The PR strategy has a minimal impact on features, and PRO-VPT demonstrates better stability than VPT in both representation and performance.
@@ 4.5 Observations on PRO-VPT
Observation 2. PRO-VPT demonstrates enhanced stability and robustness. Intuitively, runtime modifications of the prompt distribution may compromise stability. To validate this, we evaluate both the representation stability and performance stability in PRO-VPT and VPT. Specifically, to assess feature stability, we measure cosine similarity between representations (specifically [CLS] embeddings) from the previous epoch and after each stage of pruning, allocation, and tuning. This similarity reflects the extent of feature changes resulting from each stage. Fig. 6 illustrates the similarity at different blocks and the corresponding test accuracy on Cifar100. Surprisingly, the orange and green lines indicate that the PR strategy has little impact on the learned features, as the features remain nearly unchanged after pruning and allocation. Moreover, PRO-VPT exhibits superior stability compared to VPT, both in terms of representation stability and performance stability. Additionally, as shown in Fig. vi, PRO-VPT demonstrates remarkable robustness to variations in prompt quantity compared to VPT, whereas performance of VPT fluctuates significantly while PRO-VPT remains consistent. Overall, we argue that these benefits stem from our proposed optimization-based approach, which effectively identifies the optimal road for relocation, thereby minimizing its impact on stability and enhancing robustness.
Observation 3. PRO-VPT better accommodates feature hierarchies. Since different ViT blocks capture distinct hierarchical features, it is intuitive that the hierarchical nature may result in varying needs of prompts at different blocks. To validate this, we visualize the attention weights from the representation features (i.e., [CLS]) to the prompts across different blocks on Cifar100. Each attention weight reflects the importance of a prompt in relation to its corresponding feature hierarchy [38, 32]. As depicted in Fig. 7, the attention weights in VPT confirm that the hierarchical structure indeed leads to diverse prompt requirements across blocks. PRO-VPT effectively meets these block-specific needs,
Figure 7: Visualization of attention weights from features to prompts.
facilitating a more effective learning of hierarchical features.
@@ 5 Conclusion
Figure ii: Convergence curve across different prompt-tuning epochs, which corresponds to Fig. i.
@@
  • § A presents detailed explanations of the related technical concepts.
  • § B contains the complete results of distribution adjustments and offers an in-depth analysis of the underlying nature of ADO, which motivates the nested optimization formulation.
  • § C presents attempts at adding-based adjustments, which demonstrate significant instability.
  • § D explains why our proposed PR strategy relocates only a single prompt for each iteration.
  • § E details the derivation of the Taylor expansion for the idleness score.
  • § F provides more details of implementation.
  • § G provides complete results for VTAB-1k and FGVC and presents additional experimental results.
  • § H includes additional visualizations and analyses.
  • § I discusses the limitations of PRO-VPT and points out the potential direction for our future work.
@@ Appendix A Detailed Explanations of Technical Concepts
Vision Transformer. Given an input image $\mathbf{x}$, ViT [7, 24] first divides it into $n_{e}$ fixed-sized patches. Each patch is then embedded into $d$-dimensional latent space and combined with position encoding. The resulting set of patch embeddings is denoted as $\mathbf{E}_{0}\!=\!\{\mathbf{e}_{0j}\!\in\!\mathbb{R}^{d}\}^{n_{e}}_{j=1}$. A learnable classification token $\mathbf{x}_{0}$ is then concatenated with these embeddings, forming the input sequence $[\mathbf{x}_{0},\mathbf{E}_{0}]$. This sequence is then fed into a series of $L$ Transformer blocks $\{B_{i}\}^{L}_{i=1}$ as follows:
$[\mathbf{x}_{i},\mathbf{E}_{i}]=B_{i}\big{(}[\mathbf{x}_{i-1},\mathbf{E}_{i-1}]\big{)},\leavevmode\nobreak\ i=1,2,\ldots,L.$ (i)
Visual Prompt Tuning. Given a set of trainable prompt tokens $\mathbf{P}$ and the prompted backbone $f_{\mathbf{P}}$, the overall objective of VPT is to optimize these prompts for effectively adapting the PVM to downstream tasks:
$\min_{\mathbf{P}}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\mathcal{L}(f_{\mathbf{P}}(\mathbf{x}),y)\big{]}.\vspace{-0.2em}$ (ii)
This formulation specifically corresponds to Eq. (2). Depending on how the prompt set $\mathbf{P}$ is distributed across the Transformer blocks, the standard VPT [20] can be categorized into two variants, VPT-Shallow and VPT-Deep:
@@
$[\mathbf{x}_{1},\mathbf{Z}_{1},\mathbf{E}_{1}]\!$ $=\!B_{1}\big{(}[\mathbf{x}_{0},\mathbf{P},\mathbf{E}_{0}]\big{)},$ (iii)
$[\mathbf{x}_{i},\mathbf{Z}_{i},\mathbf{E}_{i}]\!$ $=\!B_{i}\big{(}[\mathbf{x}_{i-1},\mathbf{Z}_{i-1},\mathbf{E}_{i-1}]\big{)},\leavevmode\nobreak\ i\!=\!2,3,\ldots,L.$ (iv)
@@
$[\mathbf{x}_{i},\_,\mathbf{E}_{i}]=B_{i}\big{(}[\mathbf{x}_{i-1},\mathbf{P}_{i-1},\mathbf{E}_{i-1}]\big{)},\leavevmode\nobreak\ i=1,2,\ldots,L.$ (v)
@@
where ‘__\__’ indicates that VPT-Deep does not preserve the output corresponding to the prompt token $\mathbf{P}_{i-1}$.
@@
Proximal Policy Optimization. PPO is an on-policy RL algorithm that can be applied to both discrete and continuous action spaces. PPO-Clip updates policies via:
$\theta_{t+1}=\mathop{\arg\min}\limits_{\theta}\leavevmode\nobreak\ \mathbb{E}_{s,a\sim\pi_{\theta_{t}}}\big{[}L(s,a,\theta_{t},\theta)\big{]},$ (vi)

### v2

@@ 3.3 Iterative Prompt Relocation-based VPT
where the idleness score $\widehat{\mathcal{I}}_{k^{\!\ast}}\!\approx\!\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P},\mathcal{D}^{-}})$ provides an approximation of the intermediate loss of $f_{\mathbf{P},\mathcal{D}^{-}}$.
To this end, by leveraging RL to maximize the estimated reward as in Eq. (8), we can efficiently determine the optimal allocation to the prompt-needed block $a^{\ast}$ even within the non-stationary environment, thereby fulfilling Eq. (4) and forming the relocated distribution $\mathcal{D}^{+}$. Particularly, we accomplish this by employing proximal policy optimization (PPO) [48], a state-of-the-art RL algorithm.
Remark. Although the relocation process is decomposed into two-step sequential steps, the overall optimization objective, whether based on the expected objectives in Eqs. (3) and (4) or the estimated objectives in Eqs. (6) and (8), can be expressed as $\max_{k,a}\mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}[\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P}^{\prime},\mathcal{D}\mid d_{k}\!=\!a})]$. This formulation remains equivalent to the original PR objective (i.e., the ADO objective) defined in Eq. (1). Consequently, leveraging the nested optimization formulation, the proposed PR strategy can effectively address the ADO problem and adaptively construct the optimal distribution.
@@ 3.4 Discussion
We note that some prior works have indirectly achieved adaptive distribution [13, 27, 67]. Specifically, pruning-based techniques [13, 27] introduce mask variables to iteratively prune redundant prompts from the original distribution, and NOAH [67] uses an evolutionary algorithm to search for an optimal sub-distribution within the original distribution in a one-shot manner. PRO-VPT is distinct and advantageous compared to these methods as:
Distinct Goals. These methods aim to enhance performance through the lottery ticket hypothesis [8, 41, 47], focusing on searching for a more effective sub-distribution within a given hyper-distribution by circumventing negative prompts and retaining essential ones. However, they are limited to pruning prompts from prompt-saturated blocks and fail to tackle prompt-deficient blocks (cf. § 4.5). In contrast, our approach actively learns an optimal distribution of all prompts, adaptively relocating them from prompt-saturated blocks to prompt-needed blocks through our PR strategy, thereby maximizing the effectiveness of the entire prompt distribution. This fundamental distinction sets our approach apart from the methods discussed above.
Efficient. While effective, we argue that these approaches typically incur considerable unnecessary training costs to identify an optimal sub-distribution, as numerous trained prompts are pruned and left unused (cf. § 4.3). This fundamentally contradicts the goal of parameter efficiency. Conversely, our work directly adjusts the distribution of the entire prompt set through prompt relocation, thereby minimizing extra costs and improving efficiency.
Effective. These methods straightforwardly apply one-shot or iterative processes without considering the underlying distribution dynamics. For instance, NOAH [67] adopts a one-shot approach post-training, leading to suboptimal results (cf. § 4.2). Conversely, our work conducts empirical analyses and accordingly proposes a nested optimization-based method, enabling prompt-based methods to reach state-of-the-art performance.
@@ 4.1 Experiment Setups
Datasets. We evaluated our proposed PRO-VPT on a series of datasets grouped into three categories: (1) VTAB-1k. The visual task adaptation benchmark (VTAB) [66] encompasses 19 classification tasks categorized into three group: Natural, Specialized, and Structured, aiming to assess adaptation capabilities under low-shot training conditions. (2) FGVC. Fine-grained visual classification (FGVC) includes 5 classification tasks: CUB-200-2011 [56], NABirds [55], Oxford Flowers [42], Stanford Dogs [6], and Stanford Cars [10], aiming to assess task adaptation in large-scale, fine-grained settings. (3) COCO and ADE20k. These widely-used benchmarks for detection and segmentation are employed to assess the generalizability of our method across diverse downstream visual tasks.
Implementation Details. We mainly use ViT-B/16 [7] supervised pre-trained on ImageNet-21k [46] as the initialization. We employ the SGD optimizer with a batch size of 64 and fine-tune models for 100 epochs. The number of prompt tokens per task in our method is set in accordance with [22]. Following [51], we evaluate our method using Inception normalization instead of ImageNet normalization, and for fair comparisons, we present the results of other counterparts using both normalization techniques. Notably, unlike most approaches [68, 9, 22, 34], our method employs a fixed learning rate instead of using a learning rate schedule, as we necessitate ensuring timely rewinding after prompt relocation [44]. More details about the hyper-parameters are provided in Appendices G.
@@ 4.2 Main Results
VTAB-1k. We evaluate the proposed PRO-VPT on VTAB-1k and compare it with other methods. We provide results for full fine-tuning and linear probing as baselines of traditional fine-tuning. As competing advanced adapter- or reparameterization-based methods, we include SSF [34], SPT-Adapter [16], Adapter+ [51], LoRA [20], FacT-TK [23], and Consolidator [15]. Furthermore, we benchmark against state-of-the-art prompt-based methods including VPT [22], NOAH [67], SPT [58], and iVPT [71].
Tab. 1 presents results on VTAB-1k with the best-performing normalization techniques (i.e., ImageNet or Inception) for each method (see Appendix H for complete results). Among all prompt-based methods, the previous best performance of 77.1% average accuracy achieved by iVPT falls short of the advanced adapter-based method Adapter+, which reaches 77.6%. Nevertheless, with the introduction of the optimization-based adaptive distribution, the proposed PRO-VPT sets a new state-of-the-art performance among all evaluated methods with an average accuracy of 78.0%, highlighting the significant potential of prompt-based approaches. Additionally, compared to the one-shot optimization-based NOAH, our nested optimization-based PRO-VPT achieves an average accuracy gain of 2.5 pp, validating the effectiveness of our proposed nested optimization framework.
@@
Furthermore, Fig. 4 illustrates the performance differences between VPT w/ ADO (PRO-VPT) and w/o ADO (VPT-Deep) on the VTAB-1k datasets. The accuracies of PRO-VPT are consistently higher than those of VPT-Deep, yielding an average accuracy improvement of 1.6 pp. This emphasizes the significance of optimizing the prompt distribution to unlock the full potential of VPT.
@@
FGVC. Next, we evaluate PRO-VPT on the FGVC benchmark. Among the aforementioned contenders, only a portion have conducted evaluations on FGVC and released the code and configurations.
Tab. 2 presents the FGVC results with the best-performing normalization (see Appendix H for complete results). PRO-VPT achieves the best average accuracy of 91.7% over all five FGVC datasets, surpassing the second-best performance of 91.4% achieved by Adapter+ and SPT-Deep. This demonstrates that PRO-VPT also delivers state-of-the-art results for task adaptation when training data is large-scale.
@@ 4.3 Ablation Study
Strategies (A), (B) represent alternatives for extracting a sub-distribution from a hyper-distribution (see § 3.4). We utilize the pruning-based technique as in [13], applying pruning percentages ranging from 10% to 90% with 10 pp intervals. (B) prunes from a huge hyper-distribution to obtain a sub-distribution with the same final parameter count as the VPT baseline, and (A) prunes from a smaller hyper-distribution to match the number of trained parameters in the VPT baseline.
Strategies (C)-(G) represent alternatives of prompt relocating, each incorporating different components. Particularly, (C) frames the PR process as a naive RL problem (see § 3.3) and utilizes the PPO algorithm. (D)-(G) are various combinations of the following components:
@@
Observation 3. Both the pruning and allocation components in PRO-VPT exhibit significantly effective, especially RL-based allocation. Comparing strategies (D) vs. (F) and (E) vs. (G), Prn2 indicates superior performance improvements (averaging 1.07 pp for SVHN and 0.79 pp for Resisc45), highlighting the effectiveness of the proposed idleness score. Moreover, strategies (G) and (E) achieve the best results, with Alc2 delivering remarkable gains over Alc1 (2.24 pp and 1.71 pp), highlighting the advantage of framing the allocation task as a non-stationary RL problem.
@@ 4.4 Generalizability Analysis
Pre-Training Strategy. In addition to the supervised pre-training strategy, we experiment with two self-supervised strategies: the masked image modeling method (MAE) [18] and the contrastive self-supervised method (MoCo-v3) [3]. As illustrated in Tab. 5, PRO-VPT consistently outperforms VPT regardless of the pre-training strategy, highlighting its strong generalizability.
Detection and Segmentation. We also experiment on detection and segmentation tasks in Appendix H, where PRO-VPT still maintains advantages over VPT.
@@ 4.5 Observations on PRO-VPT
Observation 2. PRO-VPT effectively meets the block-specific prompt requirements via prompt relocation. We visualize the attention weights from representation features (i.e., [CLS] embeddings) to prompts across different blocks on Cifar100. Each attention weight represents the importance of a given prompt for its corresponding block. As depicted in Fig. 6, the attention weights in VPT reveal phenomena of prompt saturation (e.g., blocks 2 and 5) and prompt deficiency (e.g., blocks 4 and 12), confirming that different blocks exhibit significantly varying prompt demands. However, due to its fixed prompt distribution, VPT lacks the flexibility to adapt to these diverse needs. Even pruning-based methods can only prune negative prompts from prompt-
Figure 6: Visualization of prompt attention weights.
saturated blocks while failing to address prompt-deficient blocks. In contrast, PRO-VPT effectively meets these block-specific prompt demands through adaptive prompt relocation. Such a fine-grained tuning strategy maximizes the effectiveness of the prompt distribution, thereby enhancing downstream task performance.
Figure 7: Cosine similarity of feature changes and the corresponding test accuracy. The PR strategy has a minimal impact on features, and PRO-VPT demonstrates better stability than VPT in both representation and performance.
Observation 3. PRO-VPT demonstrates enhanced stability and robustness. One might be skeptical that runtime modifications of the prompt distribution could compromise stability. To investigate this, we evaluate both the representation stability and performance stability in PRO-VPT and VPT. Specific to feature stability, we measure cosine similarity between representations (i.e., [CLS] embeddings) from the previous epoch and after each stage of pruning, allocation, and tuning. This similarity reflects the extent of feature changes resulting from each stage. Fig. 7 illustrates the similarity at different blocks and the corresponding test accuracy on Cifar100. Surprisingly, the orange and green lines indicate that the PR strategy has little impact on the learned features, as the features remain nearly unchanged after pruning and allocation. Moreover, PRO-VPT exhibits superior stability compared to VPT, both in terms of representation stability and performance stability. Additionally, as shown in Fig. vi, PRO-VPT demonstrates remarkable robustness to variations in prompt quantity compared to VPT, whereas the performance of VPT fluctuates significantly while PRO-VPT remains consistent. Overall, we argue that these benefits stem from our proposed optimization-based approach, which effectively identifies the optimal road for relocation, thereby minimizing its impact on stability and enhancing robustness.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants 62376233, 62306181, 62476063, and 62306313; in part by the NSFC / Research Grants Council (RGC) Joint Research Scheme under Grant N_HKBU214/21; in part by the General Research Fund of RGC under the Grant 12202622; in part by the RGC Senior Research Fellow Scheme under Grant SRFS2324-2S02; in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515010163; in part by the National Key Laboratory of Radar Signal Processing under Grant JKW202403; and in part by Xiaomi Young Talents Program.
@@ Acknowledgements
  • § A provides detailed explanations of the related technical concepts.
  • § B provides the detailed pseudo-code of PRO-VPT.
  • § C presents the complete results of distribution adjustments and offers an in-depth analysis of the underlying nature of ADO, which motivates the nested optimization formulation.
  • § D presents attempts at adding-based adjustments, which demonstrate significant instability.
  • § E explains why our proposed PR strategy relocates only a single prompt for each iteration.
  • § F details the derivation of the Taylor expansion for the idleness score.
  • § G provides more details of implementation.
  • § H presents complete results for VTAB-1k and FGVC as well as additional experimental results.
  • § I presents additional visualizations and analyses.
  • § J discusses the limitations of PRO-VPT and points out the potential direction for our future work.
@@ Appendix A Detailed Explanations of Technical Concepts
Vision Transformer. Given an input image $\mathbf{x}$, ViT [7, 26, 64] first divides it into $n_{e}$ fixed-sized patches. Each patch is then embedded into $d$-dimensional latent space and combined with position encoding. The resulting set of patch embeddings is denoted as $\mathbf{E}_{0}\!=\!\{\mathbf{e}_{0j}\!\in\!\mathbb{R}^{d}\}^{n_{e}}_{j=1}$. A learnable classification token $\mathbf{x}_{0}$ is then concatenated with these embeddings, forming the input sequence $\mathbf{x}_{0},\mathbf{E}_{0}]$. This sequence is then fed into a series of $L$ Transformer blocks $\{B_{i}(\cdot)\}^{L}_{i=1}$ as follows:
$\mathbf{x}_{i},\mathbf{E}_{i}]=B_{i}\big([\mathbf{x}_{i-1},\mathbf{E}_{i-1}]\big),\penalty 10000\ i=1,2,\ldots,L.$ (i)
Visual Prompt Tuning. Given a set of trainable prompt tokens $\mathbf{P}$ and the prompted backbone $f_{\mathbf{P}}(\cdot)$, the overall objective of VPT [22] is to optimize these prompts for effectively adapting the PVM to downstream tasks:
$\min_{\mathbf{P}}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\mathcal{L}(f_{\mathbf{P}}(\mathbf{x}),y)\big].$ (ii)
This formulation specifically corresponds to Eq. (2).
Depending on how the prompt set $\mathbf{P}$ is distributed across the Transformer blocks, the standard VPT [22] can be categorized into two variants, VPT-Shallow and VPT-Deep:
@@
$\hskip-10.0pt[\mathbf{x}_{1},\mathbf{Z}_{1},\mathbf{E}_{1}]\!$ $=\!B_{1}\big([\mathbf{x}_{0},\mathbf{P},\mathbf{E}_{0}]\big),$ (iii)
$\hskip-10.0pt[\mathbf{x}_{i},\mathbf{Z}_{i},\mathbf{E}_{i}]\!$ $=\!B_{i}\big([\mathbf{x}_{i-1},\mathbf{Z}_{i-1},\mathbf{E}_{i-1}]\big),\penalty 10000\ i\!=\!2,3,\ldots,L.$ (iv)
@@
$\hskip-4.0pt[\mathbf{x}_{i},\_,\mathbf{E}_{i}]=B_{i}\big([\mathbf{x}_{i-1},\mathbf{P}_{i-1},\mathbf{E}_{i-1}]\big),\penalty 10000\ i=1,2,\ldots,L.\hskip-3.0pt$ (v)
@@
where ‘_\_’ indicates that VPT-Deep does not preserve the output corresponding to the prompt tokens $\mathbf{P}_{i-1}$.
@@
Proximal Policy Optimization. PPO [48] is a widely used RL algorithm that can be applied to both discrete and continuous action spaces. Specifically, PPO-Clip updates policies by solving the following optimization problem:
$\theta_{k+1}=\mathop{\arg\min}\limits_{\theta}\penalty 10000\ \mathbb{E}_{s,a\sim\pi_{\theta_{k}}}\big[L(s,a,\theta_{k},\theta)\big],$ (vi)